OpenStudy (anonymous):

please help!!! Simplify, showing your steps. Tell the property you used in each step. (will give a medal to first person) 1/4 (1x4) + (-4 +4)

OpenStudy (fibonaccichick666):

ok so first I will _Property 1__ \[1/4\] to the \[(1*4)\] Which property did I use?

OpenStudy (fibonaccichick666):

PS this yields \[1/4\]

OpenStudy (anonymous):

is it multiplicative property?

OpenStudy (fibonaccichick666):

ok I'm sorry I only just realized that this is not algebra but numeric only please allow me to restart: Do you know the acronym: Please Excuse My Dear Aunt Sally?

OpenStudy (anonymous):

yeah this right ? please= parenthesis excuse= exponent my= multiply dear= divide aunt= add sally=subtract

OpenStudy (fibonaccichick666):

good And that is your order of operations so let's start, what do you think is your first step?

OpenStudy (anonymous):

i look at the parenthesis first

OpenStudy (fibonaccichick666):

good let's start with the easy one (-4+4) Do you know what this equals?

OpenStudy (anonymous):

yeah zero :)

OpenStudy (fibonaccichick666):

good!! ( you would not believe how many people cannot do that ) now which property did you use?

OpenStudy (fibonaccichick666):

if you have a list it would be a good idea to look at it

OpenStudy (anonymous):

ummmm..not sure & . i dont have a list so is it additive property ?

OpenStudy (anonymous):

is it supposed to be this

OpenStudy (anonymous):

yeah that @Tahmina

OpenStudy (anonymous):

wait no the x^4 part no its just 1 times 4

OpenStudy (fibonaccichick666):

well let us first list all of our properties-Just so I know how in depth I need to go what class is this for?

OpenStudy (anonymous):

algebra 1

OpenStudy (fibonaccichick666):

college or hs?

OpenStudy (anonymous):

high school. i have geometry but im doing online classes of algebra 1 long story..

OpenStudy (fibonaccichick666):

ok that helps me so the axioms(or properties) of a field which is what you are doing the terminology may be a little advanced but it could be useful so they are: \[ a+-a=0 [Additive-inverse]\]\[a+0=a[Additive-Identity]\]\[1*a=a [Multiplicative-Identity]\]\[\frac{ 1 }{ a}*a=1[Multiplicative-Inverse]\]

OpenStudy (fibonaccichick666):

now we want to apply these properties to our problem

OpenStudy (fibonaccichick666):

let us number them 1-4 to make our lives easier

OpenStudy (fibonaccichick666):

so what form is (-4+4)?

OpenStudy (anonymous):

additive- inverse?

OpenStudy (fibonaccichick666):

good! now what does that leave us with?

OpenStudy (anonymous):

is it this ?